OpenStudy (anonymous):

How to find the inverse of 1 over (x+2)^2

OpenStudy (anonymous):

\[y = \frac{ 1 }{ \sqrt{x} }-2\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

do you want me to show you how?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

step one: \[x = \frac{ 1 }{(y+2)^2 }\]

OpenStudy (anonymous):

step two: take the sqrt of both sides \[\sqrt{x}(y+2)=1\]

OpenStudy (anonymous):

did get this step?

OpenStudy (anonymous):

\[y+2 = \frac{ 1 }{ \sqrt{x} }\]

OpenStudy (anonymous):

oh ok. so to get rid of the square, you root it

OpenStudy (anonymous):

\[y = \frac{ 1 }{ \sqrt{x} }-2\]

OpenStudy (anonymous):

\[f^{-1}(x)=\frac{ 1 }{ \sqrt{x}}-2\]

OpenStudy (anonymous):

I hope this help! ask me if you are having a problem with any step...

OpenStudy (anonymous):

when yo took the root, y did you have the y+2 beside it. Bc if I do it, I will have root x on one side and then 1 over y+2 on the other.

OpenStudy (anonymous):

yes you can do that, i just combined two steps in one

OpenStudy (anonymous):

ok, so after that what would I do? Would I have to multiply both sides by y+2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then I would divide by the root x and subtract2 from both sides

OpenStudy (anonymous):

exactly well done

OpenStudy (anonymous):

did you notice that i swapped the x and y in the first step

OpenStudy (anonymous):

yes. I remember being told to do that when I was taught about inverses. Could I not swap them and have the inverse = x

OpenStudy (anonymous):

it is easier to swap, otherwise more likely to make an error...

OpenStudy (anonymous):

oh ok. Thank you

OpenStudy (anonymous):

you are welcome! all the best...

OpenStudy (whpalmer4):

That's only half of the inverse relation, however. You didn't take the negative root into account. \[f(x) = y = \frac{1}{(x+2)^2}\]Swap variables \[x = \frac{1}{(y+2)^2}\] Take reciprocal of both sides\[\frac{1}x = (y+2)^2\]Take the square root of both sides\[\pm\frac{1}{\sqrt{x}} = (y+2)\]Subtract 2 from both sides\[f^{-1}(x)= -2\pm\frac{1}{\sqrt{x}}\] In the attached graph, you can see what happens if you only use one root: you only get the green or red portion, not both.